# InterMT: Multi-Turn Interleaved Preference Alignment with Human Feedback

### v1

@@ 1 Introduction
Humans perceive the world through dynamic, multimodal interactions involving text, images, audio, video, and more [1, 2, 3]. Building on the success multimodal large language models (MLLMs) [4, 5, 6, 7, 8], recent efforts aim to develop general-purpose AI assistants that handle multiple mixed modalities [9, 10, 11]. A key feature of such general-purpose assistants is to engage in natural multi-turn conversations, perceive and generate any modality (i.e., interleaved multimodal understanding and generation), to enable more smooth interaction and grounded understanding [9, 12, 11, 13, 14].
Recent years have seen community efforts in transplanting alignment techniques (e.g., Reinforcement Learning from Human Feedback (RLHF)) from the text modality [15, 8, 16, 17] to multiple modalities settings [13, 18, 19, 20, 21, 22, 14]. Within this line of research, most studies focus exclusively on either understanding [23, 18] or generation [22, 21]. The lack of alignment considerations for multimodal mixed input-output settings exacerbates the imbalance across modalities, i.e., modality disequilibrium.[14]. Furthermore, existing methods primarily focus on single-turn interactions, where an LLM generates a response from a prompt and receives immediate alignment feedback. However, real-world interactions typically occur in long-horizon conversations (e.g., over 5 turns) and often feature interleaved multimodal inputs and outputs [24, 25, 26].
@@
  • Modeling Long-Horizon, Interleaved Multimodal Interactions. Real-world user–AI exchanges typically span many turns and interleave text, vision, and other modalities. Such interactions demand not only precise instruction following but also sustained attention and reasoning over an evolving context, approaching near-human in-context reasoning capabilities.
  • Dynamic Human-in-the-Loop Alignment. In extended, multimodal interactions, user preferences continually evolve. For example, a user may first ask the assistant to draw a vase, then—after inspecting the rendered image—request that the vase be repositioned or restyled for greater emphasis. Capturing and aligning with these emergent, dynamic preferences calls for genuine, iterative human feedback throughout the interaction.
@@
  • One More Thing We introduce InterMT-Bench to evaluate the ability of MLLMs in assisting judges across multi-turn, multimodal tasks, encompassing three parts: Scoring Evaluation, Pair Comparison, and Crucial Step Recognition (Section 4.2). Despite strong reasoning capabilities, advanced MLLMs (e.g., o4-mini [27]) fail to align with human values in judgment tasks. However, they show potential in identifying crucial steps in long-context scenarios.
@@ 2.1 Dataset Composition
The InterMT dataset includes: (1) carefully crafted seed questions for multi-turn, multimodal conversations, and (2) fine-grained human preference annotations at both local and global conversation levels. Inspired by theories from linguistics, human-computer interaction, and cognitive psychology [28, 29, 30, 31, 32], the seed questions are rigorously selected and refined to enable more faithful simulation of real-world interleaved multimodal understanding and generation and multi-turn tasks. We collect preference data through score evaluations and pairwise comparisons of multi-modal responses at each conversation turn, based on four sub-dimensions. Global conversation helpfulness is then evaluated via five sub-dimensions. Incorporating natural language feedback further improves annotation quality and alignment with human intent. The Data Card for InterMT is as follow:
@@ 2.2 Multi-turn QA Construction
Figure 3: Overview of the four-stage pipeline for constructing InterMT. Stage I: seed questions are harvested from open-source corpora, websites, and human writing, then filtered for perceptual quality, diversity, and multi-turn potential. Stage II: iterative calls to large models and external tools (e.g. edit, retrieve, generate) produce answer expansions and follow-up questions, forming a candidate QA tree. Stage III: human annotators perform per-turn (local) and conversation-level (global) evaluations—covering quality, coherence, context awareness, and completeness—to prune and select preferred branches. Stage IV: the retained branches are reassembled into deep, coherent QA trees (depth $\geq$ 5) yielding the final multi-turn QA pairs for model training.
Prompt Collection.  InterMT is constructed from 100k image-text QA instances collected from three primary sources: 72.1% from public datasets [14, 23, 33]; 22.8% from legally scraped web content; and the remaining 5.1% from researcher-curated, human-written prompts. These instances span diverse vision-language tasks, e.g., activity generation, data visualization, and table analysis.
Drawing upon cognitive psychology theories [28, 29, 30, 31, 32], we identify five common scenarios that give rise to multi-turn conversations in real-world multimodal settings. Based on these scenarios, we filter, diversify, and rewrite the original image-text QA instances, resulting in 15604 unique seed questions. These questions serve as the initial round for generating multi-turn conversation data. Additional details can be found in Appendix D.
Tool-Augmented Agent Workflow for QA Construction.  We identify two core challenges in constructing multi-turn QA instances that capture realistic scenarios of multimodal understanding and generation: (1) How to effectively simulate realistic human multi-turn conversations in multimodal contexts? (2) Given that current MLLMs lack interleaved understanding and generation capabilities [34, 24], how to construct interleaved QA instances that generalize across diverse real-world tasks?
@@
Agent Construction.  The agent workflow is built upon a combination of strong open-source models [35, 36, 37, 4] alongside leading API-based models [38, 39, 40]. To support diverse multimodal operations, three types of image-centric tools are integrated: (1) text-to-image generators (e.g., FLUX.1-Schnell [41] and Stable-Diffusion [42]) for producing high-quality images based on prompts; (2) an image editing API (e.g., Gemini-2.0-flash [43]) capable of cropping, highlighting, and modifying images; and (3) web-based retrieval interfaces for sourcing real-world visuals. During multi-turn QA generation, agents embed structured tokens such as <Image,caption> within the text to denote visual references after which GPT-4o [38] serves as a double classifier and verifier, automatically determining the appropriate tool call based on the image intent and context.
@@ 3 Analysis
Figure 5: Agreement rates between different roles on local and global levels.
@@
Human Feedback vs. AI Feedback. Human-labeled data introduce high cost, which motivates the exploration of MLLMs’ potential to assist with evaluation tasks [44]. We develop a pipeline that utilizes advanced API-based models [38, 45, 46, 39, 47, 27]) to produce multidimensional scores from both global and local perspectives. Then, we evaluate the agreement between AI and human annotators, as well as between AI annotators and expert human verifiers. Agreement scores are then averaged across all pairs for comparative analysis. Experimental results (Fig. 5) show that while AI annotators achieve approximately 60% agreement on local evaluations, their consensus with humans on global (longer-horizon) tasks is markedly lower. This indicates current MLLMs struggle to match human judgments in multi-turn, multimodal scoring. Until AI feedback efficacy is firmly established, replacing human annotation remains inadvisable.
@@ 4.1 Preference Modeling for Multi-turn Interleaved Multimodal Scenarios
A widely adopted approach for modeling human preferences is to employ a preference predictor grounded in the Bradley–Terry (BT) model [48]. However, when extending to multi-turn settings, new challenges arise—particularly in capturing the dynamics of evolving user preferences across turns. Moreover, traditional outcome-level reward signals often fail to generalize in purely textual domains [49], let alone in complex multimodal settings involving interleaved understanding and generation. InterMT incorporates both local and global human annotations in multi-turn, multimodal interactions, leading us to investigate efficient preference modeling methods.
Inspired by [50, 51], we investigate two strategies for modeling long-horizon preferences in multi-turn multimodal scenarios: prefix preference and chain-based preference. Details of formulations can be seen in Appendix H. Our findings, presented in Figure 6, suggest that modeling fine-grained local (turn-level) preferences is more effective in capturing human values and achieving better alignment. In contrast, directly modeling global (conversation-level) preferences often fails to reflect these nuanced preferences, especially in complex, long-horizon scenarios.
@@
The results reveal two key insights: (1) Multi-turn judge moderation exhibits a generalization effect linked to the number of turns. As shown in Figure 7 (a), for evaluation turn $k$, as the number of preceding turns increases from 1 to $k-1$, the model’s accuracy continues to improve, with average future performance rising, indicating that training on multi-turn data with a limited number of turns can generalize to longer horizons. (2) Regarding the number of turns in the training data, the generalization effect shows a diminishing trend. As demonstrated in Figure 7 (b), training with $k$ turns does improve performance for $k+1\rightarrow T$ turns, but this effect diminishes as the number of turns increases. The decline is due to three factors: diminishing returns as the model struggles with long-term preferences, contextual drift as earlier turns lose relevance, and the evolving interaction between user intentions and latent preferences.
@@ 4.2 MLLM as a Judge and InterMT-Bench
Do MLLMs truly understand what is desirable in multi-turn, multimodal interactions and how to align with human values? This task is particularly challenging due to the absence of multimodal benchmarks that capture human preferences in multi-turn settings. Inspired by [52, 44] and leveraging genuine feedback from InterMT, we introduce InterMT-Bench to assess MLLMs’ alignment with human values in multi-turn, multimodal tasks. InterMT-Bench comprises three distinct tasks: Scoring Evaluation, Pair Comparison, and Crucial Step Recognition. Details can be found in Appendix G.
@@
Table 1: Overall performance comparison of different MLLMs in three judgment tasks of InterMT-Bench. All reported pearson similarity values exhibit a $p$-value below 0.05, indicating a statistically significance confidence level.
@@
Induced Bias and Hallucination.  Consistent with [44], we identify issues related to bias and hallucination: Position Bias, where models consistently favor responses in specific positions (e.g., the first answer), often influenced by training data that places correct answers at the beginning or end of prompts [53], and High-Score Bias [44], where models tend to assign higher scores to entire multi-turn communications. These issues, particularly in long-horizon tasks, may hinder the model’s ability to capture differences between extended conversations, thereby posing challenges in modeling long-horizon human values and potentially leading to safety concerns [54].
@@ 5 Conclusion and Outlook
We employed 30 full-time, experienced crowdsourced workers for multimodal annotation, ensuring fair and transparent compensation significantly above the local minimum wage, in compliance with labor regulations. The InterMT project underwent thorough ethical review by the Institutional Review Board and the Academic Committee of the Institution for Artificial Intelligence at Peking University. The dataset will be released under a CC BY 4.0 license. To mitigate potential risks, NSFW filtering was applied during dataset construction, though we acknowledge that absolute safety cannot be ensured. Given the heightened societal risks associated with multimodal data, we recommend safeguards such as gated access. We are committed to the responsible development of AI and strongly oppose any malicious or unethical use of the InterMT dataset.

### v2

@@ Abstract
Equal contribution, Corresponding author. Project website: https://pku-intermt.github.io.
@@ 1 Introduction
Humans perceive the world through dynamic, multimodal interactions involving text, images, audio, video, and more [liu2024improved, turk2014multimodal, zhang2023llama, zhang2025safevla]. Building on the success multimodal large language models (MLLMs) [liu2024llava, wu2024nextgpt, zhao2023survey, yin2023survey, ji2023ai], recent efforts aim to develop general-purpose AI assistants that handle multiple mixed modalities [alayrac2022flamingo, liu2023visual, team2024chameleon]. A key feature of such general-purpose assistants is to engage in natural multi-turn conversations, perceive and generate any modality (i.e., interleaved multimodal understanding and generation), to enable more smooth interaction and grounded understanding [alayrac2022flamingo, liu2024visual, team2024chameleon, yu2024rlhf, ji2024align, min2023factscore].
Recent years have seen community efforts in transplanting alignment techniques (e.g., Reinforcement Learning from Human Feedback (RLHF)) from the text modality [bai2022training, ji2023ai, ji2024aligner, rafailov2024direct, zhou2025sequence, ji-etal-2025-language-models, meng2025med] to multiple modalities settings [yu2024rlhf, yu2024rlaif, sun2023aligning, Maaz2023VideoChatGPT, majumder2024tango, wallace2024diffusion, ji2024align, zhou2025generative, ji2025safe]. Within this line of research, most studies focus exclusively on either understanding [zhang2025mm, yu2024rlaif] or generation [wallace2024diffusion, majumder2024tango]. The lack of alignment considerations for multimodal mixed input-output settings exacerbates the imbalance across modalities, i.e., modality disequilibrium.[ji2024align]. Furthermore, existing methods primarily focus on single-turn interactions, where an LLM generates a response from a prompt and receives immediate alignment feedback. However, real-world interactions typically occur in long-horizon conversations (e.g., over 5 turns) and often feature interleaved multimodal inputs and outputs [liu2024holistic, he2024multi, sirdeshmukh2025multichallenge].
@@
  • Modeling Long‐Horizon, Interleaved Multimodal Interactions. Real‐world user–AI exchanges typically span many turns and interleave text, vision, and other modalities. Such interactions demand not only precise instruction following but also sustained attention and reasoning over an evolving context, approaching near‐human in‐context reasoning capabilities.
  • Dynamic Human‐in‐the‐Loop Alignment. In extended, multimodal interactions, user preferences continually evolve. For example, a user may first ask the assistant to draw a vase, then—after inspecting the rendered image—request that the vase be repositioned or restyled for greater emphasis. Capturing and aligning with these emergent, dynamic preferences calls for genuine, iterative human feedback throughout the interaction.
@@
  • One More Thing We introduce InterMT-Bench to evaluate the ability of MLLMs in assisting judges across multi-turn, multimodal tasks, encompassing three parts: Scoring Evaluation, Pair Comparison, and Crucial Step Recognition (Section 4.2). Despite strong reasoning capabilities, advanced MLLMs (e.g., o4-mini [openai2025o4mini]) fail to align with human values in judgment tasks. However, they show potential in identifying crucial steps in long-context scenarios.
@@ 2.1 Dataset Composition
The InterMT dataset includes: (1) carefully crafted seed questions for multi-turn, multimodal conversations, and (2) fine-grained human preference annotations at both local and global conversation levels. Inspired by theories from linguistics, human-computer interaction, and cognitive psychology [grice1975logic, grosz1995centering, clark1991grounding, parsing2009speech, traum1995computational, pku2025deception], the seed questions are rigorously selected and refined to enable more faithful simulation of real-world interleaved multimodal understanding and generation and multi-turn tasks. We collect preference data through score evaluations and pairwise comparisons of multi-modal responses at each conversation turn, based on four sub-dimensions. Global conversation helpfulness is then evaluated via five sub-dimensions. Incorporating natural language feedback further improves annotation quality and alignment with human intent. The Data Card for InterMT is as follow:
@@ 2.2 Multi-turn QA Construction
Figure 3: Overview of the four‐stage pipeline for constructing InterMT. Stage I: seed questions are harvested from open‐source corpora, websites, and human writing, then filtered for perceptual quality, diversity, and multi‐turn potential. Stage II: iterative calls to large models and external tools (e.g. edit, retrieve, generate) produce answer expansions and follow-up questions, forming a candidate QA tree. Stage III: human annotators perform per‐turn (local) and conversation-level (global) evaluations—covering quality, coherence, context awareness, and completeness—to prune and select preferred branches. Stage IV: the retained branches are reassembled into deep, coherent QA trees (depth $\geq$ 5) yielding the final multi‐turn QA pairs for model training.
Prompt Collection.  InterMT is constructed from 100k image-text QA instances collected from three primary sources: 72.1% from public datasets [ji2024align, zhang2025mm, chen2024comm]; 22.8% from legally scraped web content; and the remaining 5.1% from researcher-curated, human-written prompts. These instances span diverse vision-language tasks, e.g., activity generation, data visualization, and table analysis.
Drawing upon cognitive psychology theories [grice1975logic, grosz1995centering, clark1991grounding, parsing2009speech, traum1995computational], we identify five common scenarios that give rise to multi-turn conversations in real-world multimodal settings. Based on these scenarios, we filter, diversify, and rewrite the original image-text QA instances, resulting in 15604 unique seed questions. These questions serve as the initial round for generating multi-turn conversation data. Additional details can be found in Appendix D.
Tool-Augmented Agent Workflow for QA Construction.  We identify two core challenges in constructing multi-turn QA instances that capture realistic scenarios of multimodal understanding and generation: (1) How to effectively simulate realistic human multi-turn conversations in multimodal contexts? (2) Given that current MLLMs lack interleaved understanding and generation capabilities [xia2025mmie, liu2024holistic], how to construct interleaved QA instances that generalize across diverse real-world tasks?
@@
Agent Construction.  The agent workflow is built upon a combination of strong open-source models [Qwen2VL, bai2025qwen2, team2025gemma, liu2024llava] alongside leading API-based models [openai2024gpt4o, google2025gemini2flash, anthropic2024claude3]. To support diverse multimodal operations, three types of image-centric tools are integrated: (1) text-to-image generators (e.g., FLUX.1-Schnell [black2024flux] and Stable-Diffusion [rombach2021highresolution]) for producing high-quality images based on prompts; (2) an image editing API (e.g., Gemini-2.0-flash [google2025geminiimagegeneration]) capable of cropping, highlighting, and modifying images; and (3) web-based retrieval interfaces for sourcing real-world visuals. During multi-turn QA generation, agents embed structured tokens such as <Image,caption> within the text to denote visual references after which GPT-4o [openai2024gpt4o] serves as a double classifier and verifier, automatically determining the appropriate tool call based on the image intent and context.
@@ 3 Analysis
Figure 5: Agreement rates between different roles at local and global levels. Here, Human refers to the outsourced annotation team, while Verifier denotes human experts with professional backgrounds and the authors, who perform random sampling and manually verify 5–10% of the dataset.
@@
Human Feedback vs. AI Feedback. Human-labeled data introduce high cost, which motivates the exploration of MLLMs’ potential to assist with evaluation tasks [chen2024mllm]. We develop a pipeline that utilizes advanced API-based models [openai2024gpt4o, openai2025gpt41, google2025geminipro, google2025gemini2flash, anthropic2025claude37, openai2025o4mini]) to produce multidimensional scores from both global and local perspectives. Then, we evaluate the agreement between AI and human annotators, as well as between AI annotators and expert human verifiers. Agreement scores are then averaged across all pairs for comparative analysis. Experimental results (Figure 5) show that while AI annotators achieve approximately 60% agreement on local evaluations, their consensus with humans on global (longer‐horizon) tasks is markedly lower. This indicates current MLLMs struggle to match human judgments in multi‐turn, multimodal scoring. Until AI feedback efficacy is firmly established, replacing human annotation remains inadvisable.
@@ 4.1 Preference Modeling for Multi-turn Interleaved Multimodal Scenarios
A widely adopted approach for modeling human preferences is to employ a preference predictor grounded in the Bradley–Terry (BT) model [bradley1952rank]. However, when extending to multi-turn settings, new challenges arise—particularly in capturing the dynamics of evolving user preferences across turns. Moreover, traditional outcome-level reward signals often fail to generalize in purely textual domains [shani2024multi], let alone in complex multimodal settings involving interleaved understanding and generation. InterMT incorporates both local and global human annotations in multi-turn, multimodal interactions, leading us to investigate efficient preference modeling methods.
Inspired by [qiu2024reward, liao2025tpo, zhang2012human], we investigate two strategies for modeling long-horizon preferences in multi-turn multimodal scenarios: prefix preference and chain-based preference. Details of formulations can be seen in Appendix H. Our findings, presented in Figure 6, suggest that modeling fine-grained local (turn-level) preferences is more effective in capturing human values and achieving better alignment. In contrast, directly modeling global (conversation-level) preferences often fails to reflect these nuanced preferences, especially in complex, long-horizon scenarios.
@@
The results reveal two key insights: (1) Multi-turn judge moderation exhibits a generalization effect linked to the number of turns. As shown in Figure 7 (a), for evaluation turn $k$, as the number of preceding turns increases from 1 to k1k-1, the model’s accuracy continues to improve, with average future performance rising, indicating that training on multi-turn data with a limited number of turns can generalize to longer horizons. (2) Regarding the number of turns in the training data, the generalization effect shows a diminishing trend. As demonstrated in Figure 7 (b), training with $k$ turns does improve performance for $k+1\rightarrow T$ turns, but this effect diminishes as the number of turns increases. The decline is due to three factors: diminishing returns as the model struggles with long-term preferences, contextual drift as earlier turns lose relevance, and the evolving interaction between user intentions and latent preferences.
@@ 4.2 MLLM as a Judge and InterMT-Bench
Do MLLMs truly understand what is desirable in multi-turn, multimodal interactions and how to align with human values? This task is particularly challenging due to the absence of multimodal benchmarks that capture human preferences in multi-turn settings. Inspired by [zheng2023judging, chen2024mllm] and leveraging genuine feedback from InterMT, we introduce InterMT-Bench to assess MLLMs’ alignment with human values in multi-turn, multimodal tasks. InterMT-Bench comprises three distinct tasks: Scoring Evaluation, Pair Comparison, and Crucial Step Recognition. Details can be found in Appendix G.
Judge Settings and Metrics
The dataset includes multi-turn multimodal interleaved communication histories and human-annotated ground truth. Evaluated models must assess the conversation at both the turn and conversation levels across nine dimensions, following a set of guidelines. Scoring Evaluation requires the model to assign scores on a 0-3 scale, with evaluation based on Pearson similarity [lee1988thirteen, zheng2023judging, chen2024mllm]. Pair Comparison directly compares two individual turns or entire conversations, without considering ties, and is evaluated for accuracy against human judgments. Crucial Step Recognition test if MLLMs can accurately identify the user’s intent and determine whether it has been fulfilled, evaluated by the score provided by judge according to the human-annotated reference answers.
@@
Table 1: Overall performance comparison of various MLLMs across three judgment tasks in InterMT-Bench. The Score Evaluation task requires models to assign scores on a 0–3 scale and is evaluated using Pearson correlation metric. The Pair Comparison task involves direct binary comparison of two individual turns or entire conversations without ties and is evaluated by accuracy against human judgments. All reported Pearson correlation values achieve statistical significance with p-values below 0.05. Note: Gemini-Flash* denotes Gemini-2.0-Flash, Gemini-Pro* corresponds to Gemini-2.5-Pro-preview, and Claude-thinking* refers to the Claude-3.7-Sonnet (thinking) model.
@@
Induced Bias and Hallucination.  Consistent with [chen2024mllm], we identify issues related to bias and hallucination: Position Bias, where models consistently favor responses in specific positions (e.g., the first answer), often influenced by training data that places correct answers at the beginning or end of prompts [li2023generative], and High-Score Bias [chen2024mllm], where models tend to assign higher scores to entire multi-turn communications. These issues, particularly in long-horizon tasks, may hinder the model’s ability to capture differences between extended conversations, thereby posing challenges in modeling long-horizon values and potentially leading to safety concerns [anil2024many].
Reasoning Ability Is Not a Panacea. We compare weak reasoning (i.e., providing plausible explanations for the evaluation results) with strong reasoning (i.e., using advanced reasoning models like o4-mini) on the scoring evaluation and pair comparison settings. However, the results are suboptimal. The models’ reasoning processes are primarily based on a step-by-step comparison against predefined guidelines, rather than actively identifying potential flaws in the responses. This approach, which differs from the more granular feedback humans provide, leads to misalignments with human preferred judgement.
Divide and Conquer is beneficial for Crucial Step Recognition We observe that models with high scores in Crucial Step Recognition (shown in Figure 9) tend to adopt a divide-and-conquer approach, meaning they first assess each turn in a multi-turn, multimodal dialogue for problem-solving, and then provide an overall conclusion; in contrast, models with lower scores often give more generalized and ambiguous responses.
@@ 5 Conclusion and Outlook
We employed 30 full-time, experienced crowdsourced workers for multimodal annotation, ensuring fair and transparent compensation significantly above the local minimum wage, in compliance with labor regulations. The InterMT project underwent thorough ethical review by the Institutional Review Board and the Academic Committee of the Institution for Artificial Intelligence at Peking University. The dataset will be released under a CC BY-NC 4.0 license. To mitigate potential risks, NSFW filtering was applied during dataset construction, though we acknowledge that absolute safety cannot be ensured. Given the heightened societal risks associated with multimodal data, we recommend safeguards such as gated access. We are committed to the responsible development of AI and strongly oppose any malicious or unethical use of the InterMT dataset.